OpenStudy (anonymous):

Show that: −ln(x−x2−1−−−−−√)=ln(x+x2−1−−−−−√)

OpenStudy (anonymous):

\[-\ln (x-\sqrt{x ^{2}-1})=\ln(x+\sqrt{x ^{2}-1})\]

OpenStudy (zzr0ck3r):

I dont think this is true

OpenStudy (anonymous):

how do i prove it?

OpenStudy (zzr0ck3r):

plug in 2 on both sides

OpenStudy (zzr0ck3r):

or 1 would be easier

OpenStudy (zzr0ck3r):

err 1 is a solution

OpenStudy (zzr0ck3r):

so plug in anything besides 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

they cancel each other

OpenStudy (anonymous):

\[-\ln (2-\sqrt{3})=\ln (2+\sqrt{3})\]

OpenStudy (zzr0ck3r):

those are not the same thing

OpenStudy (zzr0ck3r):

\(-\ln (2-\sqrt{3})=\ln (2+\sqrt{3})\\\implies \ln (\frac{1}{2-\sqrt{3}})=\ln (2+\sqrt{3})\\\implies \frac{1}{2-\sqrt{3}}=2+\sqrt{3}\)

OpenStudy (zzr0ck3r):

I don't believe it. But then again it is summer and I have been out of school for a month...

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle -\ln(x-\sqrt{x^2-1})=\ln(x+\sqrt{x^2-1}) }\) \(\large\color{black}{ \displaystyle \ln\left(\frac{1}{x-\sqrt{x^2-1}}\right)=\ln(x+\sqrt{x^2-1}) }\) \(\large\color{black}{ \displaystyle \frac{1}{x-\sqrt{x^2-1}}=x+\sqrt{x^2-1} }\) \(\large\color{black}{ \displaystyle1=\left(x+\sqrt{x^2-1} \right)\left(x-\sqrt{x^2-1}\right) }\) \(\large\color{black}{ \displaystyle1=x^2-x\sqrt{x^2-1}+x\sqrt{x^2-1}-(x^2-1) }\)

OpenStudy (solomonzelman):

and then simplify that...

OpenStudy (zzr0ck3r):

Ok so why does it seem that math is broken? i.e. what i did with x=2?

OpenStudy (zzr0ck3r):

wow I am out of it

OpenStudy (zzr0ck3r):

I was just assuming 1/(2-sqrt{3}) < 1

OpenStudy (zzr0ck3r):

that cant be a thing. you cant disprove it for x=2 and prove it for x= anything...

OpenStudy (zzr0ck3r):

But it is true for x=2. I am just slow today...

OpenStudy (zzr0ck3r):

\(\frac{1}{2-\sqrt{3}}=2+\sqrt{3}\) is true

OpenStudy (jdoe0001):

@radicalgal the above process by @SolomonZelman is correct

OpenStudy (jdoe0001):

and the graph, as show above as well, shows that both graph indeed are one and the same, thus both of them are equal indeed

OpenStudy (solomonzelman):

but there are extraneous solutions somewhere when x<0.

OpenStudy (solomonzelman):

besides these extraneous solutions, everything is correct I suppose... Thanks jdoe0001

OpenStudy (jdoe0001):

just to expand some ---> \(\bf 1=\left(x+\sqrt{x^2-1} \right)\left(x-\sqrt{x^2-1}\right) \implies 1=(x)^2-(\sqrt{x^2-1})^2 \\ \quad \\ 1=x^2-(x^2-1)\implies 1=\cancel{x^2-x^2}+1\)

OpenStudy (zzr0ck3r):

It is correct but not written correctly for a direct proof(the same way that it is not OK to manipulate both sides in a trig identity). But that depends on the level of proof expected....